OpenStudy (anonymous):

Help :/ Derivative find dy/dx y=x^5-0.125x^2+0.25x

OpenStudy (nincompoop):

the sum of a derivative is the derivative of the sum? laughing out loud

OpenStudy (nincompoop):

just apply power rule on all of them

OpenStudy (nincompoop):

do you know the power rule?

OpenStudy (anonymous):

5x^4 -.25X+0.25 right

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

\[\frac{ dy }{ dx }y=\frac{ d }{ dx }x^5-\frac{ d }{ dx }0.125x^2+\frac{ d }{ dx }0.25x\]

OpenStudy (nincompoop):

\[\frac{ d }{ dx }ax^n=a \times nx^{n-1}\]

OpenStudy (anonymous):

who can i work it out

OpenStudy (anonymous):

i did but i did my work right or not ?

OpenStudy (jdoe0001):

shorouq91 may the POWER be with you young skywalker

OpenStudy (nincompoop):

yes

OpenStudy (anonymous):

5x^4-.25x+.25

OpenStudy (nincompoop):

but you should really put y'= ... ...

OpenStudy (nincompoop):

the "prime" or dy/dx is an indication that your answer is the derivative of

OpenStudy (anonymous):

oky

OpenStudy (nincompoop):

good job

OpenStudy (anonymous):

thank you for your time